7. A heat engine has one-third the thermal efficiency of a Carnot engine operating between temperatures of 90 °C and 410 °C. If the heat engine absorbs heat at a rate of 54.0 kW, at what rate is heat exhausted?
Answer #1

T = Lower temperature = 90+273 = 363k TH= Higher temperature = 410+273 = 683k Thermal efficiency of heat engine (14.8.) - 1x
Answer #2
Chup kar ziada bakwas na kar
source: Student
answered by: Sabrina